OpenStudy (anonymous):

negative 5 over 6 as a fraction divided by negative 15 over 6 as a fraction?? show work!       

OpenStudy (aroub):

-5/6 / -15/6 -5/6 *6 / -15/6*6 -5/-15 1/3

OpenStudy (anonymous):

invert -15/6 and multiply

jhonyy9 (jhonyy9):

-5 --- 6 -5 6 1 ---- = ---- * ----- = --- -15 6 -15 3 ---- 6

OpenStudy (anonymous):

Whenever you have a fraction divided by another fraction, like jimmy said, you can "flip" the fraction and multiply: \[\frac{\frac{a}{b}}{\frac{c}{d}} = \frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

and we all love multiplying fractions because you can just multiply straight across: \[\frac{a}{b}*\frac{d}{c} = \frac{ad}{bc}\]

OpenStudy (anonymous):

Dividing is really multiplying by the multiplicative inverse (this is why you can "flip" the fraction): \[\left(-\frac{5}{6}\right) \div \left(-\frac{15}{6}\right) = \left(-\frac{5}{6}\right) \cdot \left(-\frac{15}{6}\right)^{-1} = \left(-\frac{5}{6}\right) \cdot \left(-\frac{6}{15}\right) = \frac{1}{3} \]